A sorce,S of sound,having a natural frequency n,first approches ,and then moves fast,a stationary observer O.The source moves with a constant velocity Vs of sound,in air, is V.
Find the magnitude of change delta n in the apparent frequency of the source,at the instant it goes past O.

Grade:11

2 Answers

Shivansh
5 years ago
The frequency of approach will be

f1= nV/(V-Vs)
f2= nV/(V+Vs)

Difference will be-

f1-f2 = 2nV*Vs/(V2-Vs2)
Shivansh
5 years ago
f2 as written above is the frequency heard after the source moves past